OpenStudy (anonymous):

Find the GCD for 5767 and 4453

OpenStudy (anonymous):

You mean GCF?

OpenStudy (anonymous):

The problem says GCD @LearningIsAwesome

OpenStudy (zzr0ck3r):

same thing divisor/factor

ganeshie8 (ganeshie8):

familair with euclid's algorithm ? \[\large \gcd(a, b) = \gcd(b, r)\]

ganeshie8 (ganeshie8):

\(\large a=qb+r\) *

OpenStudy (zzr0ck3r):

5767=4453*1+1314 4453=1314*3+511 1314=511*2+292 511=292*1+219 282=219*1+73 219=73*3+0 once we get to zero the last row has our GCF of 73